Use induction to prove that 0–0 4j3 = n4 + 2n3 + n2 where n > 0.
Answer #1

solution 2,410 nt & 2n3a na 20 Base case: 20 insi s 4 jo = 400)3 = 4X0 =0 RHS s nt+ an3+ n2 = 0+0+020 LAS IRMS =0 . It is tru

substitua È 4 from eance to eqn (2) =0 RA 3 3 » 73 19 :. 341 Rad 2 K ² + R² & 4(R+1) R 4 + 2k3 t R² + 4 (R3 4 1 4 3R² + 3k) =

Please give an upvote if you liked my solution. Thank you :)